Consider the following short-run production function (where L = variable input, Q = output):

Q=6L−0.4L2

Suppose that output can be sold for $10 per unit. Also assume that the firm can obtain as much of the variable input (L) as it needs at $20 per unit.

What is the marginal revenue product function (MRPL)?

$20

$60−$8L

$10

$6−$0.40L

What is the marginal factor cost function (MFCLMFCL)?

$6−$0.40L

$60−$8L

$10

$20

What is the optimal value of L, given that the objective is to maximize profits?

_________(calculate the answer) units
Answer #1

Q = 6L - 0.4L2

(1) (B)

MPL = dQ/dL = 6 - 0.8L

MRPL = MPL x Output price = $10 x (6 - 0.8L) = $60 - $8L

(2) (D)

MFCL = Wage rate = $20

(3) Profit is maximized when MRPL = MFCL

60 - 8L = 20

8L = 40

L = 5